# Imaging the Femtoscopic Freeze-out Source in Heavy Ion Collisions

### v1

@@ Abstract
By combining the femtoscopic interferometry and the optical deblurring algorithm, we implement a novel method to image the source formed in HICs, while the interaction strength between the particle pair can be simultaneously determined. The spatial distribution of the emission source has been reconstructed for protons ($p$) and antiprotons ($\bar{p}$) from the respective $pp$ and $\bar{p}\bar{p}$ correlation functions in Au+Au collisions at $\sqrt{S_{\rm NN}}=200$ GeV. Within experimental uncertainties, protons and antiprotons share the same freeze-out distribution showing higher density in the center compared to the widely assumed Gaussian shape. The results evidence the matter-antimatter symmetry in coordinate space at the moment of freeze-out before the nucleons are fully randomized in the collisions.
It is accepted that matter and antimatter are created in equal amount in the first moments after the Big Bang. Yet today, we observe a universe filled with matter far exceeding antimatter. This puzzling phenomenon is related to the profound symmetry breaking in nature and remains a challenge in nuclear and particle physics. On modern accelerators, HICs create conditions akin to those in the early universe. So after the success of observing the production of antimatter [1, 2, 3], people can further investigate the spatial evolution and decoupling of the matter and antimatter on the femtoscopic scale.
The HBT intensity interferometry, invented by Hanbury-Brown and Twiss to detect the angular radius of Siris [4], has been extensively applied in nuclear physics [5, 6] to extract the temporal-spatial feature of the emission source formed in HICs by measuring the emitted particle pair correlation function (CF) at small relative momentum. In a basic picture, the CF is the result of the low-energy scattering between the particle pair emitted from the source, and carries the information of the interaction strength, defined by the scattering length $f_{0}$ and effective range $d_{0}$ between the pair, written as ($f_{0}$, $d_{0}$) in short.
Thus, simultaneous inference of the source distribution and the interaction strength becomes feasible for either particle or antiparticle pairs. Based on the Lednický-Lyuboshitz (LL) model, STAR Collaboration has verified that protons and antiprotons exhibit nearly identical interaction properties, and the extracted Gaussian source sizes for $pp$ and $\bar{p}\bar{p}$ pairs were found to be similar [7]. Specifically, within experimental uncertainties, the parameters ($f_{0}$, $d_{0}$) are equivalent for $pp$ and $\bar{p}\bar{p}$ pairs, consistent with CPT symmetry, a fundamental principle in physical laws. The Gaussian source assumption has become widely adopted in correlation function analysis, from which significant results have been obtained in the studies of particle interactions and source distributions [8, 9, 10, 11].
@@
Despite of the difficulty in solving inverse problem, the direct imaging of the source has been achieved without introducing the analytic parameterization [15, 16, 17]. More recently, the Bayessian-based deblurring method, known as Richardson-Lucy (RL) algorithm [18, 19] which was initially developed for optical imaging [20], has gained prominence in nuclear physics [21, 22, 23, 24, 25]. Particularly, Nzabahimana et al. [22] applied the RL algorithm in source imaging in HICs, highlighting the method’s effectiveness and robustness. In this letter, combining the RL deblurring algorithm and the LL model incorporating the pair interaction, we bring the realisation to solve the inverse problem of source imaging from the experimental CFs. The source distribution and ($f_{0}$, $d_{0}$) are determined simultaneously. The similarity of the source spatial distribution of $p$ and $\bar{p}$ emission has been demonstrated.
@@
where $S$ is the source function (SF) related to the relative distance $\mathbf{r}=\mathbf{r}_{1}-\mathbf{r}_{2}$ of the particle-pair in the source rest frame, and $\Psi_{\mathbf{q}}(\mathbf{r}^{\prime})$ is the wave function calculated from the relative momentum $\mathbf{q}$ and distance $\mathbf{r}^{\prime}$ in the PRF. Based on the LL model [28], the interaction strength between particle pair is characterized by ($f_{0}$, $d_{0}$) describing the strong interaction scattering amplitude in an S-wave expansion under the effective range approximation.
@@
where the response function $P(\nu|\mu)$ represents the conditional probability that a photon with the true property $\mu$ is measured as $\nu$. Although the true distribution $\mathcal{F}(\mu)$ is unknown, it can be inferred from the measured distribution $f(\nu)$ through iterative regression starting from an arbitrary initial guess. This iterative process is illustrated in Fig. 1. Starting with $t=0$, the arbitrary initial trial $\mathcal{F}^{(0)}(\mu)$ is convolved with $P(\nu|\mu)$ to predict the measurement distribution $f^{P}(\nu)$, which is then compared to the known measurement distribution $f(\nu)$. Based on the difference between them, an amplification factor $A^{(t)}$ is calculated at each $t$th iteration. The factor $A^{(t)}$ is functioning to adjust the values of $\mathcal{F}^{(t)}(\mu)$ for the next iteration. Through the iterative process, the updated distribution $\mathcal{F}^{(t+1)}(\mu)$ progressively converges toward the true distribution $\mathcal{F}(\mu)$. The convergence criteria is invoked if $f^{P}(\nu)$ becomes sufficiently close to $f(\nu)$.
From the mathematical similarity between the KP formula (1) and the optical deblurring equation (2), it is naturally suggested that the RL algorithm can be applied iteratively to solve the SF. By evaluating the converged $\chi^{2}$ value, one can identify the optimal ($f_{0}$, $d_{0}$) for the wave function that best describes the CF, thereby achieving both SF imaging and extraction of strong interaction parameters simultaneously. For the convenience of numerical implementation, the KP formula is discretized by dividing $C$ into $N$ bins and $S$ into $M$ bins ($N\geq M$). The CF $C$ is expressed by $S$ through an $N\times M$ transformation matrix $K$ as $C_{i}=\sum_{j=1}^{M}K_{ij}S_{j}$. This transformation matrix is calculated by
@@
where $\mathbf{Q}$ denotes the total momentum of the particle pair in the source rest frame. Here, $i$ and $j$ label the bin indexes of $C$ and $S$ , respectively. In one dimensional imaging, the wave function is assumed to be uniform in the polar $\theta$ and azimuthal $\varphi$ directions, leaving only the $r$-direction to consider. Consequently, the integration over $\theta$ and $\varphi$ can be performed using Monte Carlo sampling uniformly within the spherical shell. This approach allows for iterative reconstruction of the source distribution from the discretized experimental measurements of $C$ (see Methods for a detailed description of the iterative process).
@@
The imaging of the source and the extraction of interaction parameters ($f_{0}$,$d_{0}$) can now be achieved simultaneously using this method. To test the algorithm, we start with a CF calculated from the given source distribution and known $pp$ interactions with $f_{0}=7.8$ fm and $d_{0}=2.77$ fm for $pp$ interaction. In order to mimic the real measurement, the fluctuation is introduced to each point according to the experimental uncertainty. Fig. 2 (a) shows the theoretic CF (thick solid curve) and the fluctuated CF (open circle). Then the inverse problem is iteratively solved from a initial trial of the SF, which is taken as a Gaussian distribution with $\mu=0$ center location and arbitrary deviation. The reconstructed SF ends close to the known one, as marked SF1 in  2 (b). In order to check the robustness of algorithm against the variation of the initial SF, we also test the second initial trial, where the center of the initial Gaussian is arbitrarily set at $\mu=3$ fm instead of zero peaking. With the iteration going on, the output SF gradually approaches to the known one, as illustrated by the dashed curves in Fig.  2 (b). The marked SF2 sits on top of SF1, evidencing the robustness of the iterative procedure. Fig. 2 (c) presents the closeness quantity $\chi^{2}$ over iteration times in the imaging process of the central CF. Here $\chi^{2}$ is defined as $\chi^{2}=\sum_{i}^{N}\left(\frac{C_{i}^{P}-C_{i}}{\sigma_{i}}\right)^{2}$, where $C_{i}^{P}$ represents the value of the $i^{\rm th}$ point. During the procedure, ($f_{0}$,$d_{0}$) parameters are optimized by minimum $\chi^{2}$ and the closeness between the output and the known CF is checked in every iteration till the convergence condition is reached, namely the change of $\chi^{2}$ between two successive iterations is less than $1\times 10^{-4}$ , which functions as a threshold to avoid noise amplification. Fig. 2 (d) shows the minimum $\chi^{2}$ distribution on the ($f_{0}$,$d_{0}$) plane. The optimized ($f_{0}$,$d_{0}$), marked by red cross, converge to the input one (yellow star). The consistency between the input values and the reconstructed ones in panel (a), (b) and (d) demonstrates the validity of the deblurring algorithm.
@@
Now we apply the method to the $pp$ and $\bar{p}\bar{p}$ CFs in Au+Au at center-of-mass energy of $200~{}\rm{GeV}$ per nucleon pair [29], taken by STAR collaboration [7, 30] at the Relativistic Heavy Ion Collider (RHIC) [31, 32]. The experimental CFs are inclusive and contain the contributions from the residual CFs of $p\Lambda$ and $\Lambda\Lambda$. The residual contribution, which shall be subtracted, is derived by the difference between the two fitting curves with and without the feedback corrections shown in [29]. The pure $pp$ and $\bar{p}\bar{p}$ CFs are presented in Fig. 3 (a) by the symbols, where the uncertainties are taken same as in the experiment for each data point.
@@
In the calculation of the Kernel $K_{ij}$, one needs the momentum distribution of $p$ and $\bar{p}$. The transverse momentum $p_{\rm t}$ distributions are taken from the midrapidity range in the same colliding system [33] measured by PHENIX experiment [34, 35]. The longitudinal rapidity $y_{\rm l}$ as well as the azimuth in transverse direction are uniformly distributed. The cuts $0.4<p_{\rm t}<2.5\rm{GeV}/c$ and $|y_{\rm l}|<0.7$ are taken as same as STAR data where the CFs are measured. As well noticed, the LL model provides relatively weak constraints on $d_{0}$ when analyzing experimental data. So the widely accepted theoretical and experimental approximation $d_{0}=2.8\rm~{}fm$ is fixed here for both $pp$ and $\bar{p}\bar{p}$ pairs. The convergence criterion is set to $1\times 10^{-2}$. The corresponding optimal values of $f_{0}$ obtained are as follows:
$f_{0(pp)}=(7.7\pm 0.2)\text{fm},\quad f_{0(\bar{p}\bar{p})}=(7.9\pm 0.3)\text{fm}.$ (4)
@@
The $f_{0}$ values for $pp$ and $\bar{p}\bar{p}$ pairs are identical and consistent with both the theoretical calculations and the experimental results from STAR collaboration [7].
Figure 3: (Color Online) (a) The reconstructed CFs of $pp$ (red solid curve) and $\bar{p}\bar{p}$ (blue dashed curve) pairs in comparison to the experimental CFs (red dots for $pp$ and blue circles for $\bar{p}\bar{p}$ pairs). (b) Reconstructed SFs for $p$ (red dots) and $\bar{p}$ (blue circles) in Au+Au at center of mass energy 200 GeV. The symbols are the results of direct imaging without Gaussian assumption. The SFs with Gaussian shape obtained in [7] are shown by the curves for reference.
@@
The converged CFs of $pp$ and $\bar{p}\bar{p}$ reproduce fairly the experimental data, as demonstrated by the solid and dashed curves in Fig. 3 (a), respectively. The results of the restored source imaging are plotted in panel (b) by red dots for $p$ and blue circles for $\bar{p}$. In order to estimate the uncertainty of each imaging point, numerous pseudo-CFs for $pp$ and $\bar{p}\bar{p}$ are sampled, smearing the content in each individual $q$ bin by Gaussian fluctuation, for which the center (standard deviation) is taken from the experimental value (uncertainty). Then the standard deviation of the reconstructed SFs are calculated and plotted on each data point bin-by-bin as the error bar on the SF.
@@
The direct imaging result in Fig. 3 (b) has important implications. First, Within the uncertainty, protons and antiprotons share the identical source distribution at the instant of freeze-out. In addition to the momentum distributions done in [29], the direct imaging result here completes the last piece of evidence of the matter-antimatter symmetry in coordinate space as they are produced and evolve in high energy HICs. This evidence is obtained without imposing assumptions to the shape of the source. Second, the profile of the reconstructed SFs deviates from the Gaussian distribution for both protons and antiprotons. Quantitative calculation yields the probability for the SF to be Gaussian is 00 and $2.2\times 10^{-14}$ for $p$ and $\bar{p}$, respectively. For a clear comparison, the Gaussian distribution obtained in [29] are shown by the curves in panel (b), respectively. It is evident that the emission source are more populated at smaller radius ($r<5$ fm) but less distributed at larger radius ($105<r<$ 10 fm), as zoomed in in the inset. It is consistent with the picture of an extremely fast collision. Namely, protons and antiprotons, formed by the coalescence of quarks and antiquarks created inside the overlapping region, are not fully randomized in coordinate space prior to the freeze-out instant, otherwise they are expected to extend to the outer space in Gaussian distribution. It offers a support to the recent idea of probing the initial shape of the nucleus via relativistic HIC [36, 37, 38, 39, 13, 40], since the collision is so rapid that the spatial information of the colliding nuclei is partially preserved.

### v2

# Imaging the Freeze-out Source and Extracting Strong Interaction Parameters in Relativistic Heavy Ion Collisions

@@ Abstract
Combining the femtoscopic interferometry and the optical deblurring algorithm, we implement a novel method to image the source formed in heavy ion collisions (HICs), while the interaction strength between the particle pair can be determined simultaneously. By applying this method to Au+Au collisions at $\sqrt{S_{\rm NN}}=200$ GeV, the spatial distribution of the emission source has been reconstructed for protons ($p$) and antiprotons ($\bar{p}$) from the respective $pp$ and $\bar{p}\bar{p}$ correlation functions. Within experimental uncertainties, protons and antiprotons share the same freeze-out distribution showing higher density in the center compared to the widely assumed Gaussian shape. The results evidence the matter-antimatter symmetry in coordinate space at the moment of freeze-out before the nucleons are fully randomized in the collisions.
I Introduction
It is accepted that matter and antimatter are created in equal amount in the first moments after the Big Bang. Yet today, we observe a universe filled with matter far exceeding antimatter. This puzzling phenomenon is related to the profound symmetry breaking in nature and remains a challenge in nuclear and particle physics. On modern accelerators, heavy ion collisions (HICs) create conditions akin to those in the early universe. So after the success of observing the production of antimatter [1, 2, 3], people can further investigate the spatial evolution and decoupling of the matter and antimatter on the femtoscopic scale.
The HBT intensity interferometry, invented by Hanbury-Brown and Twiss to detect the angular radius of Siris [4], has been generalized and applied extensively in nuclear physics [5, 6] to extract the temporal-spatial feature of the emission source formed in HICs by measuring the emitted particle pair correlation function (CF) at small relative momentum. In a basic picture, the CF is the result of the low-energy scattering between the particle pair emitted from the source, and carries the information of the interaction strength, defined by the scattering length $f_{0}$ and effective range $d_{0}$ between the pair, written as ($f_{0}$, $d_{0}$) in short.
Thus, simultaneous inference of the source distribution and the interaction strength becomes feasible for either particle or antiparticle pairs. Based on the Lednický-Lyuboshitz (LL) model, STAR Collaboration has verified that protons and antiprotons exhibit nearly identical interaction properties, and the extracted source size parameters (Gaussian source) for $pp$ and $\bar{p}\bar{p}$ pairs were found to be similar [7]. Specifically, within experimental uncertainties, the parameters ($f_{0}$, $d_{0}$) are equivalent for $pp$ and $\bar{p}\bar{p}$ pairs, consistent with CPT symmetry. The Gaussian source assumption has become widely adopted in correlation function analysis, from which significant results have been obtained in the studies of particle interactions and source distributions [8, 9, 10, 11].
@@ I Introduction
Despite of the difficulty in solving inverse problem, the direct imaging of the source has been achieved without introducing the analytic parameterization [15, 16, 17]. More recently, the Bayessian-based deblurring method, known as Richardson-Lucy (RL) algorithm [18, 19] which was initially developed for optical imaging [20], has gained prominence in nuclear physics [21, 22, 23, 24, 25]. Particularly, Nzabahimana et al. [22] applied the RL algorithm in source imaging in HICs, highlighting the method’s effectiveness and robustness. In this letter, combining the RL deblurring algorithm and the LL model incorporating the pair interaction, we bring the realisation to solve the inverse problem of source imaging from the experimental CFs. The source distribution and ($f_{0}$, $d_{0}$) are determined simultaneously. The similarity of the source spatial distribution of $p$ and $\bar{p}$ formed in HICs has been demonstrated.
II Method description
@@
where $S$ is the source function (SF) related to the relative distance $\mathbf{r}=\mathbf{r}_{1}-\mathbf{r}_{2}$ of the particle-pair in the source rest frame, and $\Psi_{\mathbf{q}}(\mathbf{r}^{\prime})$ is the wave function calculated from the relative momentum $\mathbf{q}$ and distance $\mathbf{r}^{\prime}$ in the PRF.
The LL model [28] provides a framework to describe the interaction strength between particle pairs through the parameters ($f_{0}$, $d_{0}$), which characterize the S-wave scattering amplitude under the effective range approximation. The model assumes that higher-order correlations involving three or more particles are negligible, the electromagnetic effects generated by the source are insignificant, and the momentum transfer from final-state interactions is small enough to apply the smoothness approximation. Within this model, the relative wave function for $pp$ and $\bar{p}\bar{p}$ pairs in the PRF is expressed as
$\psi_{\mathbf{q}}(\mathbf{r})$ $=e^{i\delta_{c}}\sqrt{A_{c}(\eta)}[e^{-i\frac{1}{2}\mathbf{q}\cdot\mathbf{r}}F(-i\eta,1,i(\frac{1}{2}\mathbf{q}\cdot\mathbf{r}+\frac{1}{2}qr))$ (2)
$+f_{c}(q)\frac{\widetilde{G}(\frac{1}{2}qr,\eta)}{r}].$
where $F(\alpha,1,z)$ is the confluent hypergeometric function, and $\widetilde{G}(\frac{1}{2}qr,\eta)$ combines Coulomb functions $G_{0}$ and $F_{0}$,
$\widetilde{G}(\frac{1}{2}qr,\eta)=\sqrt{A_{c}(\eta)}(G_{0}(\frac{1}{2}qr,\eta)+iF_{0}(\frac{1}{2}qr,\eta)),$ (3)
where $A_{c}(\eta)=\frac{2\pi\eta}{e^{2\pi\eta}-1}$ with the definition of $\eta=2/(qa_{0})$ (the Bohr radius $a_{0}=57.5\rm fm$). The S-wave scattering amplitude $f_{c}(q)$ is expanded as
$f_{c}(q)\approx[\frac{1}{f_{0}}+\frac{1}{8}d_{0}q^{2}-q\eta h(\eta^{-1})-i\frac{1}{2}qA_{c}(\eta)]^{-1},$ (4)
with $h(x=\eta^{-1})=\sum_{n=1}^{\infty}\frac{1}{n[(nx)^{2}+1]}-C+lnx$, where $C\approx 0.5772$ is the Euler constant. For identical particles, only considering S-wave, Eq. (1) can be written as following,
$C(\mathbf{q})=\int d^{3}rW_{d}(\mathbf{q},\mathbf{r}^{\prime})S(\mathbf{r}),$ (5)
where $W_{d}(\mathbf{q},\mathbf{r}^{\prime})=\frac{1}{2}[\frac{1}{4}|\psi_{\mathbf{q}}^{s}(\mathbf{r}^{\prime})+\psi_{\mathbf{q}}^{s}(-\mathbf{r}^{\prime})|^{2}+\frac{3}{4}|\psi_{\mathbf{q}}^{t}(\mathbf{r}^{\prime})-\psi_{\mathbf{q}}^{t}(-\mathbf{r}^{\prime})|^{2}]$.
@@
where the response function $P(\nu|\mu)$ represents the conditional probability that a photon with the true property $\mu$ is measured as $\nu$. Although the true distribution $\mathcal{F}(\mu)$ is unknown, it can be inferred from the measured distribution $f(\nu)$ through iterative regression starting from an arbitrary initial guess. This iterative process is illustrated in Fig. 1. Starting with $t=0$, the arbitrary initial trial $\mathcal{F}^{(0)}(\mu)$ is convolved with $P(\nu|\mu)$ to predict the measurement distribution $f^{P}(\nu)$, which is then compared to the known measurement distribution $f(\nu)$. Based on the difference between them, an amplification factor $A^{(t)}$ is calculated at each $t^{\rm th}$ iteration. The factor $A^{(t)}$ is functioning to adjust the values of $\mathcal{F}^{(t)}(\mu)$ for the next iteration. Through the iterative process, the updated distribution $\mathcal{F}^{(t+1)}(\mu)$ progressively converges toward the true distribution $\mathcal{F}(\mu)$. The convergence criteria is invoked if $f^{P}(\nu)$ becomes sufficiently close to $f(\nu)$.
From the mathematical similarity between the KP formula (1) and the optical deblurring equation (6), it is naturally suggested that the RL algorithm can be applied iteratively to solve the SF. By evaluating the converged $\chi^{2}$ value, one can identify the optimal ($f_{0}$, $d_{0}$) for the wave function that best describes the CF, thereby achieving both SF imaging and extraction of strong interaction parameters simultaneously. For the convenience of numerical implementation, the KP formula is discretized by dividing $C$ into $N$ bins and $S$ into $M$ bins ($N\geq M$). The CF $C$ is expressed by $S$ through an $N\times M$ transformation matrix $K$ as
$C_{i}=\sum_{j=1}^{M}K_{ij}S_{j}.$ (7)
Once the values of $C$ and $K$ are known, the distribution of $S$ can be optimized through the RL algorithm via iterative processes:
$S_{i}^{(t+1)}=S_{i}^{(t)}\frac{\sum_{j}W_{j}\frac{C_{j}}{C_{j}^{(t)}}K_{ji}}{\sum_{j’}W_{j’}K_{j’i}},$ (8)
where $W$ represents the weight of the data. According to the measurement errors $\epsilon$ of each data point of CFs, the weight $W$ of $i^{\rm th}$ CF point is set to $1/\epsilon_{i}$ for the iterative process. $C^{(t)}$ is the predicted CF at the $t^{\rm th}$ iteration:
$C_{i}^{(t)}=\sum_{j}K_{ij}S_{j}^{(t)}.$ (9)
Additionally, it is crucial to consider a moderate number of iterations and an appropriate number of bins to ensure the algorithm’s robustness.
The transformation matrix $K_{ij}$ is calculated by
@@
where $\mathbf{Q}$ denotes the total momentum of the particle pair in the source rest frame. Here, $i$ and $j$ label the bin indexes of $C$ and $S$ , respectively. In one dimensional imaging, the wave function is assumed to be uniform in the polar $\theta$ and azimuthal $\varphi$ directions, leaving only the $r$-direction to consider. Consequently, the integration over $\theta$ and $\varphi$ can be performed using Monte Carlo sampling uniformly within the spherical shell. This approach allows for iterative reconstruction of the source distribution from the discretized experimental measurements of $C$.
[…]
@@
III Method test
The imaging of the source and the extraction of interaction parameters ($f_{0}$,$d_{0}$) can now be achieved simultaneously using this method. To test the algorithm, we start with a CF calculated from the given source distribution and known $pp$ interactions. The known source function is modeled as a 1D Gaussian distribution with a mean of 00 and a standard deviation of $\sigma=3.5~{}\rm fm$ in the radial direction, as illustrated by blue solid curve in Fig. 2 (b). For imaging purposes, the source function is uniformly divided into $M=20$ bins with a width of $1\rm fm$ over the range of 00 to $20~{}\rm fm$ in the radial direction, while the CF is divided into $N=50$ bins with a width of $4~{}\rm MeV$ over the range of 00 to $200~{}\rm MeV$. The transformation matrix is calculated using the LL model, incorporating Coulomb interaction, with the interaction parameters for $pp$ interactions specified as $f_{0}=7.8~{}\rm fm$ and $d_{0}=2.77~{}\rm fm$. The CF used for imaging is obtained by multiplying the transformation matrix with the source function, incorporating the corresponding error values for each point, as shown by the blue line with the purple error band in Fig. 2 (a). In order to mimic the real measurement, the fluctuation is introduced to each point according to the experimental uncertainty. By treating the value $C$ of each point in the CF as the mean of a Gaussian distribution and the associated error as its standard deviation, we can sample numerous CFs. One of these is randomly selected to serve as the central CF, represented by the symbols of hollow blue circle in Fig. 2 (a).
Each sampled CF is then iteratively solved using the imaging method. Since the exact values of ($f_{0}$,$d_{0}$), key parameters of the transformation matrix, are unknown, it is reasonable to generate different matrices by exploring all possible values of ($f_{0}$,$d_{0}$). In this study, we consider values of ($f_{0}$,$d_{0}$) within the range of 00 to $10~{}\rm fm$, with a step size of $0.1~{}\rm fm$. For each sampled CF, we solve it using these matrices and compute the corresponding $\chi^{2}$ values ($\text{NDF}=N$) which represents for the overall deviation. Here $\chi^{2}$ is defined as
$\chi^{2}=\sum_{i}^{N}\left(\frac{C_{i}^{P}-C_{i}}{\sigma_{i}}\right)^{2},$ (11)
where $C_{i}^{P}$ represents the value of the $i^{\rm th}$ point. The matrix that yields the minimum $\chi^{2}$ is considered the best fit to describe the CF, allowing us to simultaneously determine ($f_{0}$,$d_{0}$), and the imaged source function.
@@ III Method test
The initial source function for the iteration is all assumed to be a Gaussian distribution with $\sigma_{0}=3~{}\rm fm$. Then the inverse problem is iteratively solved from a initial trial of the SF, which is taken as a Gaussian distribution with $\mu=0$ center location and arbitrary deviation. The reconstructed SF ends close to the known one, as marked SF1 in  2 (b). In order to check the robustness of algorithm against the variation of the initial SF, we also test the second initial trial, where the center of the initial Gaussian is arbitrarily set at $\mu=3$ fm instead of zero peaking. With the iteration going on, the output SF gradually approaches to the known one, as illustrated by the dashed curves in Fig.  2 (b). The marked SF2 sits on top of SF1, evidencing the robustness of the iterative procedure. Fig. 2 (c) presents the closeness quantity $\chi^{2}$ over iteration times in the imaging process of the central CF. During the procedure, ($f_{0}$,$d_{0}$) parameters are optimized by minimum $\chi^{2}$ and the closeness between the output and the known CF is checked in every iteration till the convergence condition is reached, namely the change of $\chi^{2}$ between two successive iterations is less than $1\times 10^{-4}$ , which functions as a threshold to avoid noise amplification. Fig. 2 (d) shows the minimum $\chi^{2}$ distribution on the ($f_{0}$,$d_{0}$) plane. The optimized ($f_{0}$,$d_{0}$), marked by red cross, converge to the input one (yellow star). The consistency between the input values and the reconstructed ones in panel (a), (b) and (d) demonstrates the validity of the deblurring algorithm.
Once all sampled CFs have been solved, we can calculate the standard deviations of ($f_{0}$,$d_{0}$), and each bin of the imaged source function, which represent the uncertainties in the imaging results. The solution obtained from the selected central CF is considered the central value. The corresponding $\chi^{2}$ values for different matrices of the central CF are shown in Fig. 2 (d), where the red star marks the optimal ($f_{0}$,$d_{0}$). The final solution for ($f_{0}$,$d_{0}$) is
$f_{0}=(7.5\pm 0.2)\rm~{}fm,d_{0}=(2.9\pm 1.5)\rm~{}fm,$ (12)
and the imaged source function is represented by the hollow red circles in Fig. 2 (b). In Fig. 2 (a), the red solid line represent the CF reconstructed using imaging methods. In addition, due to the introduction of a $1~{}\rm fm$ cutoff when calculating the Coulomb-inclusive wave function using the LL model, in order to prevent numerical divergence during the solution process, the first point in the source function is unreliable.
IV Application to data
@@
Now we apply the method to the $pp$ and $\bar{p}\bar{p}$ CFs in Au+Au at $\sqrt{S_{\rm NN}}=200~{}\rm{GeV}$ [29], taken by STAR collaboration [7, 30] at the Relativistic Heavy Ion Collider (RHIC) [31, 32]. The experimental CFs are inclusive and contain the contributions from the residual CFs of $p\Lambda$ and $\Lambda\Lambda$. The residual contribution, which shall be subtracted, is derived by the difference between the two fitting curves with and without the feedback corrections shown in [29]. Specifically, the difference between the fitting results of inclusive CFs and $1+x_{pp(\bar{p}\bar{p})}[C_{pp(\bar{p}\bar{p})}-1]$ are subtracted based on the experimental data. The relative fraction of protons and antiprotons are estimated by the simulation of THMINATOR2 model under the global DCA cut condition (DCA$<2cm$). The values of $x_{pp}$ and $x_{\bar{p}\bar{p}}$ are 0.45 and 0.42, respectively. In addition, we suppose that there are no differences between inclusive CF and true $pp(\bar{p}\bar{p})$ CF when $q_{rel}>160\rm MeV$, which means we use the experimental data of $q_{rel}>160\rm MeV$ directly. Subsequently, by combining the values of $x_{pp(\bar{p}\bar{p})}$, the true $pp$ ($\bar{p}\bar{p}$) CFs can be obtained. The pure $pp$ and $\bar{p}\bar{p}$ CFs are presented in Fig. 4 (a) by the symbols, where the uncertainties are taken same as in the experiment for each data point.
@@
In the calculation of the Kernel $K_{ij}$, one needs the momentum distribution of $p$ and $\bar{p}$. The transverse momentum $p_{\rm t}$ distributions are taken from the midrapidity range in the same colliding system [33] measured by PHENIX experiment [34, 35]. The longitudinal rapidity $y_{\rm l}$ as well as the azimuth in transverse direction are uniformly distributed. The cuts $0.4<p_{\rm t}<2.5\rm{GeV}/c$ and $|y_{\rm l}|<0.7$ are taken as same as STAR data where the CFs are measured. And then, the three-dimension momentum distributions can be obtained. Specifically, we sample the rapidity uniformly within $-0.7<y_{1}<0.7$ ($y_{1}=\frac{1}{2}ln(\frac{1+\beta_{z}}{1-\beta_{z}})$), the $p_{t}$ from the distributions reported by the PHENIX experiment satisfied with the cuts of STAR, and the azimuthal angle $\varphi$ uniformly between 0 and $2\pi$. From these, we obtain the momentum of one particle randomly,
$p_{x}=p_{t}cos\varphi,p_{y}=p_{t}sin\varphi,p_{z}=\pm\sqrt{\frac{\beta_{z}^{2}(p_{t}^{2}+m_{p(\bar{p})}^{2})}{1-\beta_{z}^{2}}}.$ (13)
Based on the momentum sampling method described above, we present the three-dimensional momentum distributions for both protons and antiprotons under the given conditions. The momentum values required to generate the matrix are then sampled from these distributions.
As well noticed, the LL model provides relatively weak constraints on $d_{0}$ when analyzing experimental data. So the widely accepted theoretical and experimental approximation $d_{0}=2.8\rm~{}fm$ is fixed here for both $pp$ and $\bar{p}\bar{p}$ pairs. The convergence criterion is set to $1\times 10^{-2}$. The corresponding optimal values of $f_{0}$ obtained are as follows:
$f_{0(pp)}=(7.7\pm 0.2)\text{fm},\quad f_{0(\bar{p}\bar{p})}=(7.9\pm 0.3)\text{fm},$ (14)
@@
and the corresponding distributions of $f_{0}$ are shown in the Fig. 3. The $f_{0}$ values for $pp$ and $\bar{p}\bar{p}$ pairs are identical and consistent with both the theoretical calculations and the experimental results from STAR collaboration [7].
Figure 3: (Color Online) The distribution of the optimal $f_{0}$ values obtained from the imaging of each sampled pseudo-CFs. The red and blue lines represent the optimal $f_{0}$ distributions for $pp$ pairs and $\bar{p}\bar{p}$ pairs, respectively.
@@
The converged CFs of $pp$ and $\bar{p}\bar{p}$ reproduce fairly the experimental data, as demonstrated by the solid and dashed curves in Fig. 4 (a), respectively. The results of the restored source imaging are plotted in panel (b) by red dots for $p$ and blue circles for $\bar{p}$. The red and blue lines are represented for the extracted radii of $p$ and $\bar{p}$ source by STAR, respectively. The extracted radii refer to the single-particle distribution in the source rest frame, which differs from the $S(\mathbf{r})$ defined in this work. In our case, $S(\mathbf{r})$ is a function of the relative distance between two particles. To facilitate comparison, the Gaussian source radii extracted by the STAR collaboration have been multiplied by $\sqrt{2}$. In order to estimate the uncertainty of each imaging point, numerous pseudo-CFs for $pp$ and $\bar{p}\bar{p}$ are sampled, smearing the content in each individual $q$ bin by Gaussian fluctuation, for which the center (standard deviation) is taken from the experimental value (uncertainty). Then the standard deviation of the reconstructed SFs are calculated and plotted on each data point bin-by-bin as the error bar on the SF.
Figure 4: (Color Online) (a) The reconstructed CFs of $pp$ (red solid curve) and $\bar{p}\bar{p}$ (blue dashed curve) pairs in comparison to the experimental CFs [7] (red dots for $pp$ and blue circles for $\bar{p}\bar{p}$ pairs). (b) Reconstructed SFs for $p$ (red dots) and $\bar{p}$ (blue circles) in Au+Au at center of mass energy 200 GeV. The symbols are the results of direct imaging without Gaussian assumption. The SFs with Gaussian shape obtained in [7] are shown by the curves for reference.
@@
The direct imaging result in Fig. 4 (b) has important implications. First, Within the uncertainty, protons and antiprotons share the identical source distribution at the instant of freeze-out. In addition to the momentum distributions done in [29], the direct imaging result here completes the last piece of evidence of the matter-antimatter symmetry in coordinate space as they are produced and evolve in high energy HICs. This evidence is obtained without imposing assumptions to the shape of the source. Second, the profile of the reconstructed SFs deviates from the Gaussian distribution for both protons and antiprotons. To test whether the imaged source functions for $pp$ and $\bar{p}\bar{p}$ from STAR data are Gaussian distributions, each was fitted with a Gaussian function, and the corresponding chi-square (goodness of fit) values were obtained. For $pp$ and $\bar{p}\bar{p}$, the chi-square values were 91243.5 and 101.078, respectively, with 16 degrees of freedom (19 data points minus 3 parameters in the Gaussian fit). According to the chi-square test, quantitative calculation yields the probabilities for the SFs to be Gaussian is 00 and $2.2\times 10^{-14}$ for $p$ and $\bar{p}$, respectively. Thus, the Gaussian distribution hypothesis can be excluded. For a clear comparison, the Gaussian distribution obtained in [29] are shown by the curves in panel (b), respectively. It is evident that the emission source are more populated at smaller radius ($r<5$ fm) but less distributed at larger radius ($85<r<$ 8 fm), as zoomed in in the inset. It is consistent with the picture of an extremely fast collision. Namely, protons and antiprotons, formed by the coalescence of quarks and antiquarks created inside the overlapping region, are not fully randomized in coordinate space prior to the freeze-out instant, otherwise they are expected to extend to the outer space in Gaussian distribution. It offers a support to the recent idea of probing the initial shape of the nucleus via relativistic HIC [36, 37, 38, 39, 13, 40], since the collision is so rapid that the spatial information of the colliding nuclei is partially preserved.